# Measurement Error and Counterfactuals in Quantitative Trade and Spatial Models

### v4

@@ Abstract
Counterfactuals in quantitative trade and spatial models are functions of the current state of the world and the model parameters. Common practice treats the current state of the world as perfectly observed, but there is good reason to believe that it is measured with error. This paper provides tools for quantifying uncertainty about counterfactuals when the current state of the world is measured with error. I recommend an empirical Bayes approach to uncertainty quantification, and show that it is both practical and theoretically justified. I apply the proposed method to the settings in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022) and find non-trivial uncertainty about counterfactuals.
@@ 1 Introduction
This nonstandard structure is further complicated by the fact that data in quantitative trade and spatial models are often measured with error (Goes, 2023; Linsi, Burgoon, and Mügge, 2023; Teti, 2023). This gives rise to a novel measurement error problem: unlike conventional settings, where the estimand is a function of the correctly measured distribution of the data, here it depends on the correctly measured realizations. An additional complication arises when the estimand also depends on a structural parameter that is itself estimated using the noisy data. As a result, inference must account for three distinct sources of uncertainty: (i) estimation error, (ii) the direct effect of measurement error on the estimand, and (iii) the indirect effect of measurement error through the estimation of structural parameter.
To fix ideas, consider the canonical Armington model (Armington, 1969). In this model, proportional changes in welfare due to proportional changes in trade costs can be expressed as a function of baseline bilateral trade flows and the trade elasticity (Arkolakis, Costinot, and Rodríguez-Clare, 2012). Because the trade elasticity is unknown, it is typically estimated using the same trade flow data. The central question I address is how measurement error in the observed bilateral trade flows translates to uncertainty in the predicted welfare changes. As the Armington model illustrates, this uncertainty arises both directly—through the mismeasured data used in the welfare formula—and indirectly—through the estimation of the trade elasticity based on the same mismeasured data.
@@
I outline a general Bayesian framework for quantifying uncertainty that incorporates these various sources of uncertainty. The framework requires researchers to specify both a measurement error model and a prior distribution over the latent true data. For settings where the observed data consist of non-negative dyadic flows, I recommend default choices for both components that can be calibrated using the observed data, yielding a default empirical Bayes (EB) approach. Specifically, I suggest modeling measurement error as log-normal and using a log-normal prior centered on a gravity equation, with additional mass at zero to accommodate zero flows. This setup is designed for easy implementation and is well suited to a wide range of quantitative trade and spatial models.
To illustrate the impact of accounting for measurement error, I revisit the settings in Adao, Costinot, and Donaldson (2017) and Allen and Arkolakis (2022). For the counterfactual analysis in Adao, Costinot, and Donaldson (2017), which quantifies the welfare impacts of China’s accession to the WTO, I model measurement error in bilateral trade flows. I apply my default EB approach, calibrating the prior and measurement error model using the mirror trade dataset compiled by Linsi, Burgoon, and Mügge (2023). This dataset reports bilateral trade flows as recorded by both exporters and importers, which I interpret as two independent noisy measurements of the latent true trade flow. I plot the estimated change in China’s welfare from 1996 to 2011 and construct uncertainty intervals that reflect both measurement error and estimation error.
@@
In the setting of Allen and Arkolakis (2022), the counterfactual question concerns which highway links in the United States yield the highest return on investment and thus are most promising for improvement. I model measurement error in traffic flows and apply the default EB approach, calibrating the prior and measurement error model using estimates from Musunuru and Porter (2019). I compute uncertainty intervals that account for both measurement error and estimation error for the three links with the highest estimated returns. Although these intervals are wide, the relative ranking of the top three links remains robust.
@@
This paper contributes to a growing body of work aimed at improving counterfactual analysis in quantitative trade and spatial models (Balistreri and Hillberry, 2008; Adao, Costinot, and Donaldson, 2017; Kehoe, Pujolas, and Rossbach, 2017; Adão, Costinot, and Donaldson, 2023; Ansari, Donaldson, and Wiles, 2024; Sanders, 2025). The most closely related work is Dingel and Tintelnot (2020), which studies calibration procedures in granular environments. That paper considers models that presume a continuum of agents and shows that with limited data, unit-level idiosyncrasies are absorbed into the model, leading to overfitting and poor out-of-sample performance. My focus is on the complementary issue of uncertainty quantification due to measurement and estimation error—an issue that persists even in non-granular settings. Dingel and Tintelnot (2020) recommends using fitted values from a low-dimensional model in place of the raw observed data. I show how this recommendation can be naturally integrated into the Bayesian framework.
@@
The remainder of the paper is organized as follows. The next section introduces the notation and the setting I consider. Section 3 discusses how to jointly account for estimation error and measurement error in quantitative trade and spatial models. Section 4 presents my default empirical Bayes approach for uncertainty quantification. Section 5 applies the method to the trade setting in Adao, Costinot, and Donaldson (2017) and explores its use in the economic geography framework of Allen and Arkolakis (2022). Section 6 concludes.
@@ 2.1 Notation and Key Assumption
To begin, consider a baseline setting without estimation or measurement error. Let $D\in\mathcal{D}\subseteq\mathbb{R}^{d_{D}}$ denote a data vector drawn from distribution $\mathcal{P}_{D}$, and let $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$ denote a structural parameter. Our objective is to compute a scalar counterfactual quantity $\gamma\in\mathbb{R}$. The key assumption that the counterfactual object of interest has to satisfy is:
@@
Assumption 1 implies that if the data $D$ are observed without error and the structural parameter $\theta$ is known, we can perfectly recover $\gamma$ [^1] Indeed, by fixing $g$ I abstract away from model misspecification, an important problem I engage with in future work. This is different from conventional economic models, where the object of interest is a function of the correctly measured distribution of the data, rather than the actual observations. So the key distinction with conventional settings is:
$\begin{cases}\begin{array}[]{c}\mathrm{conventional:}\\ \mathrm{this\>paper:}\end{array}&\begin{array}[]{c}\gamma=g\left(\mathcal{P}_{D},\theta\right)\\ \gamma=g\left(D,\theta\right),\quad D\sim\mathcal{P}_{D}.\end{array}\end{cases}$ (2)
@@
2.1.1 Running Example: Armington Model
For illustration, I will derive the function $g$ in a simple example. A canonical workhorse model in international trade is the Armington model (Armington, 1969), as outlined, for example, in Costinot and Rodríguez-Clare (2014). Countries are indexed by $i,j=1,...,n$, and with CES preferences and perfect competition, it follows that the relevant gravity equations and budget constraints are:
$E_{i}$ $=\left(1+\kappa_{i}\right)Y_{i},\qquad i=1,...,n.$ (4)
Here, $F_{ij}$ denotes the trade flow from country $i$ to $j$, and $Y_{i}=\sum_{\ell=1}^{n}F_{i\ell}$, $E_{i}=\sum_{k=1}^{n}F_{ki}$ and $\kappa_{i}=\left(E_{i}-Y_{i}\right)/Y_{i}$ denote country $i$’s total income, total expenditure and the ratio of the trade deficit to income, respectively. Furthermore, $\tau_{ij}$ denotes the iceberg trade cost between country $i$ and $j$, which means that in order to sell one unit of a good in country $j$, country $i$ must ship $\tau_{ij}\geq 1$ units, with $\tau_{ii}=1$. Lastly, $\varepsilon>0$ is the trade elasticity and $\left\{\chi_{ij}\right\}$ are idiosyncratic terms.
Now, say we are interested in the counterfactual where we change the trade costs $\left\{\tau_{ij}\right\}$ proportionally by $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$, holding the trade elasticity $\varepsilon$, the idiosyncratic terms $\left\{\chi_{ij}\right\}$ and the trade imbalance variables $\left\{\kappa_{i}\right\}$ constant. In Appendix B.1 I show that we can then solve for the corresponding proportional changes in income, $\left\{Y_{i}^{\mathrm{cf,prop}}\right\}$, using
where $\lambda_{ij}=F_{ij}/E_{j}$ denotes the expenditure share that country $j$ spends on goods from country $i$. By Walras’ Law, the proportional changes in income are only pinned down up to a multiplicative constant. Subsequently, following Costinot and Rodríguez-Clare (2014), we can exactly solve for proportional changes in expenditure shares and welfare levels:
$\lambda_{ij}^{\mathrm{cf,prop}}$ $=\frac{\left(\tau_{ij}^{\mathrm{cf,prop}}Y_{i}^{\mathrm{cf,prop}}\right)^{-\varepsilon}}{\sum_{k}\lambda_{kj}\left(\tau_{kj}^{\mathrm{cf,prop}}Y_{k}^{\mathrm{cf,prop}}\right)^{-\varepsilon}},\qquad i,j=1,...,n,$
$D$ =$=\left\{F_{ij}\right\}$
$\theta$ $=\varepsilon.$
The specific counterfactual question I consider is a 10% increase in all bilateral trade costs between 76 countries, so that $\tau_{ij}^{\mathrm{cf,prop}}=1+0.1\cdot\mathbb{I}\left\{i\neq j\right\}$ for $i,j=1,...,n$. I focus on the changes in welfare in the Central African Republic, the Netherlands, Sweden and the United States. It follows that, fixing $\left\{\tau_{ij}^{\mathrm{cf,prop}}\right\}$, we have
$\gamma_{q}=100\cdot\left(W_{q}^{\mathrm{cf,prop}}-1\right)\equiv g_{q}\left(\left\{F_{ij}\right\},\varepsilon\right),$
for each $q\in\left\{\mathrm{CAF},\mathrm{NLD},\mathrm{SWE},\mathrm{USA}\right\}$. In Appendix B.2, I show the results for all 76 countries in the sample.
3 Bayesian Uncertainty Quantification
@@
This section introduces estimation error and measurement error to quantitative trade and spatial models. It outlines how to jointly account for them and how to quantify uncertainty for the counterfactual prediction of interest.
3.1 Introducing Estimation Error
The counterfactual prediction of interest will typically depend on a structural parameter $\theta$. In practice, $\theta$ is unknown and must be estimated from the data, yielding an estimator $\tilde{\theta}\left(D\right)$. It is common in applied work to treat this estimate as fixed—whether it is taken from the literature or obtained through data-driven methods—thus ignoring the uncertainty associated with the estimation process. An exception is Adao, Costinot, and Donaldson (2017), which reports confidence sets for the counterfactual predictions of interest that account for estimation error.
I will take a Bayesian or Quasi-Bayesian approach and assume that the posterior or quasi-posterior distribution of the true structural parameter $\theta$ given the data $D$ is approximately normal. [^2] Formally, this normality could follow from assumptions on the underlying data generating process such that a Bernstein-von Mises type result holds (Van der Vaart, 2000). In that case the influence of the prior distribution $\pi\left(\theta\right)$ becomes negligible and the posterior distribution approximately equals a normal distribution centered at the maximum likelihood estimator. See Appendix C for a discussion on justifying Equation (5) in this setting. In Sanders (2025) I engage further with structural estimation in quantitative trade and spatial models. Specifically, we have
$\pi^{EE}\left(\theta|D\right)\approx\mathcal{N}\left(\tilde{\theta}\left(D\right),\tilde{\Sigma}\left(D\right)\right),$ (5)
where $\tilde{\Sigma}\left(D\right)$ is a consistent estimator of the sampling variance of $\tilde{\theta}\left(D\right)$. [^3] This notation nests the scenario where we use an estimator from another study that used different data. In that case $\theta$ is independent from $D$ and we would write $\pi^{EE}\left(\theta|D\right)\approx\mathcal{N}\left(\tilde{\theta},\tilde{\Sigma}\right)$. Furthermore, in the case where $\theta$ is known to be non-negative, one can use a log-normal distribution here.
We can generate draws from this posterior distribution of $\theta$ given $D$. For each of these draws, we can calculate the corresponding value of the counterfactual object of interest using the relationship $\gamma=g\left(D,\theta\right)$. This allows us to find the posterior distribution of $\gamma$ given the true data, $\pi^{EE}\left(\gamma|D\right)$.
3.1.1 Running Example: Armington Model (Continued)
As we saw before, for the Armington model we have $D=\left\{F_{ij}\right\}$ and $\theta=\varepsilon$. In practice, the trade flows are also used to estimate the trade elasticity $\varepsilon$. In particular, we can rearrange the gravity equation in Equation (3) to find
$F_{ij}$ $=\exp\left\{\underbrace{-\varepsilon\log Y_{i}}_{\equiv\delta_{i}^{\mathrm{orig}}}+\underbrace{\log E_{j}-\log\sum_{k}\left(\tau_{kj}Y_{k}\right)^{-\varepsilon}\chi_{kj}}_{\equiv\delta_{j}^{\mathrm{dest}}}-\varepsilon\log\tau_{ij}+\underbrace{\log\chi_{ij}}_{\equiv\delta_{ij}}\right\}.$
If we then treat $\delta_{ij}=\log\chi_{ij}$ as a preference shock orthogonal to log trade costs $\log\tau_{ij}$, this results in the non-linear regression model
$F_{ij}=\exp\left\{\delta_{i}^{\mathrm{orig}}+\delta_{j}^{\mathrm{dest}}-\varepsilon\log\tau_{ij}+\delta_{ij}\right\}.$
In the presence of zero trade flows, this suggests using a PPML estimator as in Silva and Tenreyro (2006). To obtain a valid estimate of $\tilde{\Sigma}\left(\left\{F_{ij}\right\}\right)$ that accounts for dyadic dependence, I use results from Graham (2020).
A key challenge is that trade costs $\left\{\tau_{ij}\right\}$ are not directly observed. To address this, I use estimates from Waugh (2010), which are based on an arbitrage condition and follow the methodology of Eaton and Kortum (2002). Waugh (2010) argues that these estimates are subject to minimal measurement error. Accordingly, I treat them as accurate and proceed under the assumption that there is no measurement error in this proxy for trade costs. Appendix B.3 provides further details.
The resulting point estimate and standard error are $2.26$ and $0.52$, respectively. If we assume there is no measurement error, we can find $\pi^{EE}\left(\gamma_{q}|\left\{F_{ij}\right\}\right)$ for $q\in\left\{\mathrm{CAF},\mathrm{NLD},\mathrm{SWE},\mathrm{USA}\right\}$ by sampling trade elasticities from $\pi^{EE}\left(\varepsilon|\left\{F_{ij}\right\}\right)\approx\mathcal{N}\left(2.26,0.52^{2}\right)$ and compute the corresponding changes in welfare. This results in the point estimates and intervals as in Table 1 and the posteriors plotted in Figure 1. We observe that the Netherlands and Sweden suffer the largest welfare losses. Note that all posteriors are non-Gaussian which results in asymmetric intervals.
Point estimate Only est. error
$\gamma_{\mathrm{CAF}}$0.72 [-0.95, -0.13]
$\gamma_{\mathrm{NLD}}$5.47 [-5.50, -5.37]
$\gamma_{\mathrm{SWE}}$3.77 [-4.14, -3.54]
$\gamma_{\mathrm{USA}}$1.09 [-1.09, -1.03]
Table 1: Uncertainty quantification for the Armington model, considering only estimation error.
Figure 1: Uncertainty quantification for the Armington model, considering only estimation error.
3.2 Introducing Measurement Error
@@
Under Assumption 1, our object of interest can be written as a function solely of the data realizations and the structural parameter, which is convenient for answering counterfactual questions. However, the data realizations are economic variables which are often measured with error. For instance, Ortiz-Ospina and Beltekian (2018) and Goes (2023) highlight that there are large discrepancies between and within various data sources from trade and international economics. Motivated by this, I assume that, instead of the true data vector $D$, we observe a noisy version $\tilde{D}$. [^4] Alternatively, one can interpret the mapping from true data to observed data as a realization from a stochastic process, rather than purely as classical measurement error. The Bayesian framework I propose accommodates this alternative interpretation as well, since any such stochastic relationship between the true and observed data can be encoded via a likelihood function $\pi(\tilde{D}\mid D)$.
@@
For uncertainty quantification for the counterfactual prediction, we will require the posterior distribution of the true data given the noisy data. Towards that end, I introduce a model for the measurement error and a prior distribution for the true underlying data. Given such a prior distribution and a measurement error model,
@@
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\pi\left(D\right)\\ \pi\left(\tilde{D}|D\right),\end{array}\end{cases}$

### v5

@@ Abstract
Counterfactuals in quantitative trade and spatial models are functions of the current state of the world and the model parameters. Common practice treats the current state of the world as perfectly observed, but there is good reason to believe that it is measured with error. This paper provides tools for quantifying uncertainty about counterfactuals when the current state of the world is measured with error. I recommend an empirical Bayes approach to uncertainty quantification, and show that it is both practical and theoretically justified. I apply the proposed method to the settings in adao2017nonparametric and allen2022welfare and find non-trivial uncertainty about counterfactuals.
@@ 1 Introduction
The setting with data-dependent counterfactual estimands is further complicated by the fact that data in quantitative trade and spatial models are often measured with error (goes2023reliability; linsi2023problem; teti2023missing). Unlike classical measurement error settings, where the estimand is typically a parameter of the correctly measured population distribution, here it is a functional of the realized data. To illustrate, consider the canonical Armington model (armington1969theory), where predicted welfare changes from hypothetical trade cost shocks can be written as a function of baseline bilateral trade flows and the trade elasticity (arkolakis2012new). The question I address is how measurement error in the observed trade flows affects uncertainty in the welfare predictions.
I develop an empirical Bayes framework for quantifying uncertainty around counterfactual predictions. The approach requires specifying both a measurement error model and a prior distribution over the latent true data, up to a set of hyperparameters. These hyperparameters are estimated from the observed data via an empirical Bayes step. Bayes’ rule then yields an estimated posterior distribution over the latent data given the noisy observations. Given the structure of quantitative trade and spatial models, this posterior induces a corresponding posterior over counterfactual predictions. Uncertainty can then be summarized by reporting an interval of posterior quantiles.
A natural accompanying point estimator for this procedure is the posterior median, which is always guaranteed to be contained in the reported interval. By contrast, the standard point estimator, which does not account for measurement error, may lay outside the interval. The posterior median answers the question: What does a Bayesian believe the counterfactual prediction would have been in the absence of measurement error? While this is a natural and intuitive question to ask, the answer necessarily depends on the prior.
@@
In settings where the observed data consist of non-negative dyadic flows, I propose a default specification for the measurement error model and prior that can be calibrated directly from the data, yielding a widely applicable empirical Bayes approach. Specifically, I model measurement error as log-normal and use a log-normal prior centered on a structural gravity equation, with a point mass at zero to accommodate zero flows. This setup is designed for ease of implementation and is suitable for a wide class of quantitative trade and spatial models.
I consider two approaches to calibrating the hyperparameters in this default specification. The first assumes constant measurement error variances across flows and relies on researcher input or domain knowledge. The second, applicable in the case of international trade, uses the mirror trade dataset compiled by linsi2023problem, which reports bilateral trade flows as recorded by both exporters and importers. I interpret these paired observations as two independent noisy measurements of the true trade flow, enabling calibration of flow-specific measurement error variances.
To illustrate the impact of incorporating measurement error into counterfactual analysis, I revisit the applications in adao2017nonparametric and allen2022welfare. In adao2017nonparametric, which quantify the welfare effects of China’s accession to the WTO, I model measurement error in baseline bilateral trade flows. I apply the default empirical Bayes approach and construct uncertainty intervals that account for measurement error in the estimated changes in China’s welfare from 1996 to 2011. These intervals are substantially wider than those reported in adao2017nonparametric, which reflect estimation uncertainty.
@@
In the setting of allen2022welfare, the counterfactual question concerns which highway links in the United States yield the highest return on investment and are therefore most promising for improvement. I model measurement error in traffic flows and apply the default empirical Bayes approach, calibrating the prior and measurement error model using estimates from musunuru2019applications. I compute uncertainty intervals that account for measurement error for the three links with the highest estimated returns. Although the intervals are wide, the relative ranking of the top three links remains robust.
@@
This paper contributes to a growing body of work aimed at improving counterfactual analysis in quantitative trade and spatial models (balistreri2008gravity; adao2017nonparametric; kehoe2017quantitative; adao2023putting; ansari2024quantifying; sanders2025new). The most closely related work is DingelTintelnot:2025, which studies calibration procedures in granular environments. That paper considers models that presume a continuum of agents and shows that, when data are limited, unit-level idiosyncrasies are absorbed into the model parameters, leading to overfitting and poor out-of-sample performance. My focus is on the complementary issue of uncertainty quantification due to measurement error—an issue that persists even in non-granular settings. DingelTintelnot:2025 recommends replacing raw observed data with fitted values from a low-dimensional model. I show how this recommendation can be nested into the proposed Bayesian framework.
@@
The remainder of the paper is organized as follows. Section 2 introduces the setting and notation. Section 3 presents the empirical Bayes framework for accounting for measurement error in quantitative trade and spatial models. Section 4 describes a widely applicable default approach. Section 5 demonstrates the procedure in the context of the Armington model. Section 6 applies the method to the trade setting in adao2017nonparametric and explores its use in the economic geography framework of allen2022welfare. Section 7 concludes.
@@ 2.1 Notation and Key Assumption
To begin, consider a baseline setting where there is no measurement error. Let $D\in\mathcal{D}\subseteq\mathbb{R}^{d_{D}}$ denote a data vector drawn from distribution $\mathcal{P}_{D}$, and let $\theta\in\Theta\subseteq\mathbb{R}^{d_{\theta}}$ denote a structural parameter. Our objective is to compute a scalar counterfactual quantity $\gamma\in\mathbb{R}$. The key assumption that the counterfactual object of interest has to satisfy is:
@@
The main appeal of focusing on objects of the form in Assumption 1 is that it allows researchers to answer counterfactual questions posed relative to a specific, observed factual situation. In quantitative trade and spatial settings, such questions are often at least as relevant as those concerning average effects. For example, in a quantitative model of international trade, the goal is typically to understand what would happen to the world following a specific policy change, rather than what would occur in a randomly drawn year under that policy.
Assumption 1 implies that if the data $D$ are observed without error and the structural parameter $\theta$ is known, we can perfectly recover $\gamma$. [^1] Indeed, by fixing $g$ I abstract away from model misspecification, an important problem I engage with in future work. This contrasts with standard econometric models, where the object of interest is a function of the correctly measured distribution of the data, rather than the actual observations. So the key distinction with standard settings is:
$\begin{cases}\begin{array}[]{c}\mathrm{standard\ setting:}\\ \mathrm{this\>paper:}\end{array}&\begin{array}[]{c}\gamma=g\left(\mathcal{P}_{D},\theta\right)\\ \gamma=g\left(D,\theta\right),\quad D\sim\mathcal{P}_{D}\end{array}.\end{cases}$ (2)
@@
3 Empirical Bayes Uncertainty Quantification
@@
This section introduces measurement error into quantitative trade and spatial models. It outlines how to quantify the resulting uncertainty for the counterfactual prediction of interest.
3.1 Prior and Measurement Error Model
@@
Under Assumption 1, our object of interest can be written as a function solely of the data realizations and the structural parameter, which is convenient for answering relevant counterfactual questions. However, the data realizations are economic variables which are often measured with error. For instance, ortiz2018international and goes2023reliability highlight that there are large discrepancies between and within various data sources from trade and international economics. Motivated by this, I assume that, instead of the true data vector $D$, we observe a noisy version $\tilde{D}$.
@@
For uncertainty quantification for the counterfactual prediction, we will require the posterior distribution of the true data given the noisy data. Towards that end, I introduce a model for the measurement error and a prior distribution for the true underlying data,
@@
$\begin{cases}\begin{array}[]{c}\mathrm{prior:}\\ \mathrm{measurement\>error:}\end{array}&\begin{array}[]{c}\pi^{\mathrm{prior}}\left(D|\vartheta\right)\\ \pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)\end{array}.\end{cases}$
Here, $\vartheta\in\mathbb{R}^{d_{\vartheta}}$ is a vector of unknown hyperparameters.
3.2 Empirical Bayes and Posterior Distribution
Given such a prior distribution and a measurement error model, we can use an empirical Bayes approach to estimate the unknown hyperparameters. [^2] Rather than estimating the parameters of the prior distribution for the true underlying data, which corresponds to an empirical Bayes approach, one could alternatively specify prior distributions for these parameters, which corresponds to a hierarchical Bayes approach. Formally, we have
$\tilde{\vartheta}=\underset{\vartheta}{\arg\max}\int\pi^{\mathrm{me}}\left(\tilde{D}|D,\vartheta\right)\pi^{\mathrm{prior}}\left(D|\vartheta\right)dD.$